# S3E: A Large-scale Multimodal Dataset for Collaborative SLAM

### v5

@@ Abstract
Collaborative Simultaneous Localization and Mapping (C-SLAM), Multi-Agent Systems, Dataset, Trajectory Paradigms
@@ I Introduction
Collaborative Simultaneous Localization and Mapping (C-SLAM) is pivotal for enabling multi-agent cooperation, thereby enhancing the robustness and efficiency of concurrent localization and mapping endeavors within a shared operational space [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. Despite considerable progress in C-SLAM research during the past decade, issues of repeatability and benchmarking remain significant hurdles within the domain. The field is primarily constrained by two limitations. First, the complexity of C-SLAM systems, which integrate intricate software architectures and a variety of sensors, complicates the process of achieving perfect replication, thereby emphasizing the need for publicly available datasets to facilitate methodological evaluation. Second, a common method for evaluating C-SLAM solutions involves partitioning a single-agent SLAM dataset into several segments, each assigned to a distinct virtual agent [11]. However, this methodology is impractical for real-world multi-agent scenarios, given the uniform perspectives and lighting conditions that all virtual agents would encounter at intersection points.
@@
Within the SLAM research community, publicly available datasets are indispensable for accelerating the development of SLAM algorithms suitable for real-world deployment. This significance stems from two primary factors. First, the execution of specialized SLAM experiments requires substantial investment in sophisticated hardware and complex software, including calibration routines and ground truth generation systems; hence, the availability of datasets substantially aids in the algorithmic development process. Second, these datasets provide a fair and impartial benchmarking framework for comparative analysis and performance assessment. Over the past decade, an array of SLAM datasets has emerged, each contributing uniquely to the diversity of sensor modalities and the breadth of scenarios and motion types they encompass [12, 13, 14, 15, 16, 17]. Nonetheless, the preponderance of these datasets is skewed towards single-agent systems, equipped with an assortment of sensors such as cameras, inertial measurement units (IMUs), and LiDAR, which may limit their applicability to multi-agent contexts.
@@ III-D Trajectory Paradigms
.1 Refer to caption S3E. .2 Refer to caption <platform_id>. .3 Refer to caption Calibration. .4 Refer to caption <agent_id>.yaml. .3 Refer to caption S3E_<sequence_id>. .4 Refer to caption S3E_<sequence_id>.db3. .4 Refer to caption metadata.yaml. .4 Refer to caption groudtruth.txt.
@@ III-F Dataset Format
Data Organization: As depicted in LABEL:Fig:data_organization , our dataset features a well-organized directory structure. The calibration parameters are documented in YAML files and stored in the dedicated Calibration directory. Each sequence is accompanied by essential data files, including .db3 file that house the primary sensor measurements, and metadata.yaml file that provide essential information about the sequence, such as sensor settings and capture conditions. Furthermore, to facilitate comprehensive performance evaluation, we have included auxiliary files such as groundtruth.txt, which contains the ground truth data.

### v6

# S3E: A Mulit-Robot Multimodal Dataset for Collaborative SLAM

@@ Abstract
Note to Practitioners
The S3E dataset is a valuable asset for professionals in robotics and autonomous systems, offering a rich, multimodal set of data to refine Collaborative Simultaneous Localization and Mapping (C-SLAM) technologies. With synchronized streams from LiDAR, cameras, IMUs, and UWB across indoor and outdoor environments, it presents a diverse range of real-world challenges for testing and developing C-SLAM algorithms. This dataset is particularly beneficial for evaluating how robots perform in collaborative tasks, with its unique trajectory paradigms simulating various operational scenarios. Practitioners can use S3E to enhance their systems’ ability to handle dynamic environments, spatial constraints, and limited sensor overlap. However, working with S3E requires significant computational resources and expertise in C-SLAM. The complexity of the data may be daunting for those new to the field. Despite this, the S3E dataset is an excellent platform for pushing the boundaries of collaborative robotics, providing a rigorous benchmark for current technologies and a foundation for future developments.
@@
Collaborative Simultaneous Localization and Mapping, Multi-Agent Systems, Dataset, Trajectory Paradigms
@@ I Introduction
Collaborative Simultaneous Localization and Mapping (C-SLAM) is pivotal for enabling multi-agent cooperation, thereby enhancing the robustness and efficiency of concurrent localization and mapping endeavors within a shared operational space [1, 2, 3, 4, 5, 6, 7]. Despite considerable progress in C-SLAM research during the past decade, issues of repeatability and benchmarking remain significant hurdles within the domain. The field is primarily constrained by two limitations. First, the complexity of C-SLAM systems, which integrate intricate software architectures and a variety of sensors, complicates the process of achieving perfect replication, thereby emphasizing the need for publicly available datasets to facilitate methodological evaluation. Second, a common method for evaluating C-SLAM solutions involves partitioning a single-agent SLAM dataset into several segments, each assigned to a distinct virtual agent [8]. However, this methodology is impractical for real-world multi-agent scenarios, given the uniform perspectives and lighting conditions that all virtual agents would encounter at intersection points.
@@
Within the SLAM research community, publicly available datasets are indispensable for accelerating the development of SLAM algorithms suitable for real-world deployment. This significance stems from two primary factors. First, the execution of specialized SLAM experiments requires substantial investment in sophisticated hardware and complex software, including calibration routines and ground truth generation systems; hence, the availability of datasets substantially aids in the algorithmic development process. Second, these datasets provide a fair and impartial benchmarking framework for comparative analysis and performance assessment. Over the past decade, an array of SLAM datasets has emerged, each contributing uniquely to the diversity of sensor modalities and the breadth of scenarios and motion types they encompass [9, 10]. Nonetheless, the preponderance of these datasets is skewed towards single-agent systems, equipped with an assortment of sensors such as cameras, inertial measurement units (IMUs), and LiDAR, which may limit their applicability to multi-agent contexts.
@@ III-D Trajectory Paradigms
.1 Refer to caption S3E. .2 Refer to caption <platform_id>. .3 Refer to caption Calibration. .4 Refer to caption <agent_id>.yaml. .3 Refer to caption S3E_<sequence_id>. .4 Refer to caption S3E_<sequence_id>.db3. .4 Refer to caption metadata.yaml. .4 Refer to caption <agent_id>_gt.txt.
@@ III-F Dataset Format
Data Organization: As depicted in LABEL:Fig:data_organization , our dataset features a well-organized directory structure. The calibration parameters are documented in YAML files and stored in the dedicated Calibration directory. Each sequence is accompanied by essential data files, including .db3 file that house the primary sensor measurements, and metadata.yaml file that provide essential information about the sequence, such as sensor settings and capture conditions. Furthermore, to facilitate comprehensive performance evaluation, we have included auxiliary files such as <agent_id>_gt.txt, which contains the ground truth data.